# Variational inference via radial transport

### v1

@@ 6 Conclusion
We propose and analyze a framework for variational inference over the space of radial transport maps, leading to our algorithm, radVI. Under standard assumptions, we prove convergence guarantees for our algorithm in both the deterministic and stochastic settings. Our analysis hinges on novel regularity properties of optimal transport maps between radially symmetric distributions. We demonstrate our ability to learn radial distributions in a suite of experiments where standard VI methods fail to capture the correct behavior. An interesting open question is to lift the log-concave assumptions and still obtain optimization guarantees in this setting, as many have in the sampling literature (Balasubramanian et al., 2022; Chewi, 2026).
@@ Acknowledgements
As $\nabla^{2}(-\log\rho)=I$ and by Proposition 3.4, it holds by a two-sided version of Caffarelli’s contraction theorem (see, e.g., Chewi and Pooladian, 2023) that
@@
We use the fact that for any $\alpha$-strongly log-concave distribution $\pi$ in dimension $d$ with mode $x_{\star}$, it holds that $\int\|x-x_{\star}\|^{2}\,\mathrm{d}\pi(x)\leq d/\alpha$ (c.f. Chewi, 2026, “basic lemma”). Let $X_{\rho}\sim\rho$, so that $\Psi^{\star}(\|X_{\rho}\|)$ is distributed according to the radial part of $\pi_{\rm rad}^{\star}$. The radial parts of $\rho$ and of $\pi_{\rm rad}^{\star}$ are one-dimensional distributions, with potentials $F_{\rho}$ and $F_{\star}$ respectively; this implies that they are both strongly log-concave, with respective parameters $1$ and $\ell_{V}$. Therefore,
@@ Appendix C Omitted proofs from Section 4
We will closely follow the proof strategy put forth by Jiang et al. (2025) with appropriate modifications throughout.
@@
where $X\sim\rho$. Using standard tools from high-dimensional probability pertaining to Gaussian tail bounds (Van Handel, 2014), the right-hand side is bounded by $O(\ell_{V}^{-1}\exp(-R^{2}/2))$. Therefore, if $R\gtrsim\sqrt{\log(1/\varepsilon)}$, then $(\mathsf{T1})\leq\varepsilon^{2}/\ell_{V}$.
@@
Next, for $r\leq\sqrt{d}-R$, $\hat{\Psi}(r)=r\Psi^{\star}(\sqrt{d}-R)/(\sqrt{d}-R)$, so $\hat{\psi}(r)\leq\Psi^{\star}(\sqrt{d}-R)/(\sqrt{d}-R)\leq 1/\sqrt{\ell_{V}}$. For $r\geq\sqrt{d}+R$, $\hat{\Psi}(r)=\Psi^{\star}(\sqrt{d}+R)$, so $\hat{\psi}(r)\leq\Psi^{\star}(\sqrt{d}+R)/(\sqrt{d}+R)\leq 1/\sqrt{\ell_{V}}$. Since $\psi^{\star}\leq 1/\sqrt{\ell_{V}}$ everywhere,
$d\int_{\sqrt{d}+R}^{\infty}(\psi^{\star}-\hat{\psi})^{2}\,\mathrm{d}\widetilde{\rho}$ $\leq\frac{d}{\ell_{V}}\,\mathbb{P}(\|X\|\notin\mathcal{I})\lesssim\frac{d\exp(-R^{2}/2)}{\ell_{V}}\,.$
Hence, $R\gtrsim\sqrt{\log(d/\varepsilon)}$ implies $\int_{\|x\|\notin\mathcal{I}}\|D(T^{\star}-\hat{T})(x)\|_{\rm F}^{2}\,\mathrm{d}\rho(x)\lesssim\varepsilon/\ell_{V}$.
@@ C.2 Construction of the Gram matrix
We follow the arguments of Jiang et al. (2025). By Jiang et al. (2025, Propositions 5.10 and 5.11), the objective $\lambda\mapsto\mathcal{F}(T_{\lambda})$ is $\ell_{V}$-strongly convex and $L_{V}\,(1+\Upsilon)$ smooth in the $\|\cdot\|_{Q}$ norm, where $\Upsilon>0$ is the smallest positive constant such that $\|DT_{\lambda}-\alpha I\|_{L^{2}(\rho)}^{2}\leq\Upsilon\,\|T_{\lambda}-\alpha\,{\operatorname{id}}\|_{L^{2}(\rho)}^{2}$ for all $\lambda\in\mathbb{R}^{J+1}$. Writing this out, we need
@@
Let us start with the case of $k\geq 1$. In this case, $r\gtrsim\sqrt{d}$, so the term $\mathsf{T}_{2}$ is bounded, up to an absolute constant, by the integrand on the right-hand side. Thus, for term $\mathsf{T}_{1}$, it suffices to prove $\int_{a_{k}}^{a_{k}+\delta}\,\mathrm{d}\widetilde{\rho}(r)\lesssim\Upsilon\inf_{\bar{r}\in\mathbb{R}}\int_{a_{k}}^{a_{k}+\delta}(r-\bar{r})^{2}\,\mathrm{d}\widetilde{\rho}(r)$, i.e., we need a lower bound on the variance of the distribution $\widetilde{\rho}$ restricted to the interval $a_{k},a_{k}+\delta]$. Note that $\widetilde{\rho}\propto\exp(-F_{\rho})$ where $F_{\rho}$ is defined in the proof of Theorem 3.5. Using the estimates from that proof, we know that $F_{\rho}$ is $O(R)$-Lipschitz on the interval $a_{k},a_{k}+\delta]$, hence $\log\widetilde{\rho}$ only varies by $O(1)$ on this interval provided $\delta\lesssim 1/R$. The argument of Jiang et al. (2025, Lemma 5.13) now shows that the variance is lower bounded by $\Omega(\delta^{2})$, and hence our desired estimate holds for $\Upsilon\asymp\delta^{-2}$.
@@
Recall that for $\lambda\mapsto\mathcal{F}(T_{\lambda})$, the smoothness constant is $\widetilde{O}(J^{2}L_{V})$. Choosing $h=1/\widetilde{\Theta}(L_{V}J^{2})$, we have by standard arguments for projected gradient descent for smooth and strongly convex functions (Beck, 2017, Theorem 10.29)
@@ C.5 Gradient of the objective
$DT_{\lambda}(x)=\frac{\alpha r+\sum_{j=0}^{J}\lambda_{j}\Psi_{j}(r)}{r}I+\Big(\sum_{j=0}^{J}\lambda_{j}\Psi_{j}^{\prime}(r)-\frac{1}{r}\sum_{j=0}^{J}\lambda_{j}\Psi_{j}(r)\Big)\frac{xx^{\top}}{r^{2}}\,$,
@@
$DT_{\lambda}(x)=\Bigl(\alpha+\sum_{j=0}^{J}\lambda_{j}\Psi^{\prime}_{j}(r)\Bigr)xx^{\top}/r^{2}+\Bigl(\alpha+\sum_{j=0}^{J}\Psi_{j}(r)/r\Bigr)(I-xx^{\top}/r^{2})\,.$
@@
$\log\det DT_{\lambda}(x)=(d-1)\log\Big(\alpha+\langle\lambda,\vec{\Psi}(r)/r\rangle\Big)+\log\Big(\alpha+\langle\lambda,\vec{\Psi}^{\prime}(r)\rangle\Big),$
@@
$\nabla_{\lambda_{j}}\,\mathbb{E}_{X\sim\rho}[\log\det DT_{\lambda}(X)]=\int_{0}^{\infty}\left[\frac{(d-1)\Psi_{j}(r)}{g_{\lambda}(r)}+\frac{\Psi^{\prime}_{j}(r)}{g^{\prime}_{\lambda}(r)}\right]\,\mathrm{d}\widetilde{\rho}(r)$ (25)
@@
To prove the main result, we follow the strategy of Jiang et al. (2025, Lemma 5.16). Setting up notation, recall that
@@
where we bound $\langle Q^{-1},\vec{\Psi}(\|x\|)\vec{\Psi}(\|x\|)^{\top}\rangle\lesssim J^{3}$. This follows by mimicking Lemma 5.15 by Jiang et al. (2025) but using the computations from the proof of Proposition 4.4; we omit this computation. We bound this last term as follows
@@
The statement follows by employing Theorem 4.3 by Jiang et al. (2025).
@@
We conclude by invoking a standard fact about log-smooth measures (Chewi, 2026, “basic lemma”). ∎
@@
Our goal is to show that $Q^{-1}$ is lower-bounded by some absolute constant $c_{Q}>0$. This is sufficient, as
$\langle Q^{-1},\vec{\Psi}(x)\vec{\Psi}(x)^{\top}\rangle\frac{1}{c_{Q}}\operatorname{tr}(\vec{\Psi}(x)\vec{\Psi}(x)^{\top})=\frac{1}{c_{Q}}\sum_{j=0}^{J}\Psi_{j}^{2}(\|x\|)\leq\frac{M}{c_{Q}}\,$,
@@ D.1 Hyperparameters
To obtain a Laplace approximation of a posterior, we use the “minimize” function in Scipy (Virtanen et al., 2020) with the BFGS optimizer. Our implementation of Forward-Backward VI is exactly as in the publicly available respository by Diao et al. (2023), and we use the same learning rate and number of iterations as our approach.
@@ D.2.1 Student-$t$ distribution
Let $\pi_{\rm Stu}$ be the Student-$t$ distribution with $\nu$ degrees of freedom in $\mathbb{R}^{d}$, mean $\mu$, and covariance $\Sigma$. We say $X\sim\pi_{\rm Stu}$ if $X=\mu+Z/\sqrt{W/\nu}$ where $Z\sim\mathcal{N}(0,\Sigma)$ and $W\sim\chi^{2}_{\nu}$ are independent.
The density of a $d$-dimensional Student-$t$ distribution with degrees of freedom $\nu>0$, mean $\mu\in\mathbb{R}^{d}$, and scale matrix $\Sigma\in\mathbb{R}^{d\times d}$ is given by
$\pi_{\rm Stu}(x)=\frac{\Gamma\left(\frac{\nu+d}{2}\right)}{\Gamma\left(\frac{\nu}{2}\right)(\nu\pi)^{d/2}|\Sigma|^{1/2}}\Bigl(1+\frac{r(x)^{2}}{\nu}\Bigr)^{-\frac{\nu+d}{2}}.$
@@
In the radial setting, it suffices to enforce this on balls of radius $r$ i.e., on sets of the form $\{x:\|x\|\leq r\}$, yielding, for a random variable $X\sim\rho$ and $Y\sim\pi,$
@@
and $R$ is determined implicitly by solving
@@
If $Y\sim\pi_{\rm Stu}$, then $\|Y\|^{2}/d\sim F_{d,\nu}$, where $F_{d,\nu}$ is the CDF of the FF–distribution with $(d,\nu)$ degrees of freedom. Hence,
$\mathbb{P}(\|Y\|\leq s)=\mathbb{P}\!\left(\frac{\|Y\|^{2}}{d}\leq\frac{s^{2}}{d}\right)=F_{d,\nu}\!\left(\frac{s^{2}}{d}\right),$
@@
$F_{\chi^{2}_{d}}(r^{2})=F_{d,\nu}\!\left(\frac{\Psi^{\star}(r)^{2}}{d}\right),$
@@
$\Psi^{\star}(r)=\sqrt{\,d\,F^{-1}_{d,\nu}\!\big(F_{\chi^{2}_{d}}(r^{2})\big)}.$
@@ D.2.2 Laplace distribution
Solving the implicit equation for the mapped radius $\Psi^{\star}(r)$ point-wise is computationally prohibitive. To circumvent this, we use an interpolation-based approach: the radial CDF of the target is pre-computed on a dense grid, and a linear interpolator of its inverse is constructed. The map then reduces to a direct, vectorized evaluation of this interpolator. Alternatively, the equation can be solved for each point using a root-finding algorithm, such as Brent’s method (Brent, 1973), although this approach is less efficient for large datasets.

### v2

@@ 6 Conclusion
We propose and analyze a framework for variational inference over the space of radial transport maps, leading to our algorithm, radVI. Under standard assumptions, we prove convergence guarantees for our algorithm in both the deterministic and stochastic settings. Our analysis hinges on novel regularity properties of optimal transport maps between radially symmetric distributions. We demonstrate our ability to learn radial distributions in a suite of experiments where standard VI methods fail to capture the correct behavior. An interesting open question is to lift the log-concave assumptions and still obtain optimization guarantees in this setting, as many have in the sampling literature (Balasubramanian et al.,, 2022; Chewi,, 2026).
@@ Acknowledgements
As $\nabla^{2}(-\log\rho)=I$ and by Proposition 3.4, it holds by a two-sided version of Caffarelli’s contraction theorem (see, e.g., Chewi and Pooladian,, 2023) that
@@
We use the fact that for any $\alpha$-strongly log-concave distribution $\pi$ in dimension $d$ with mode $x_{\star}$, it holds that $\int\|x-x_{\star}\|^{2}\,\mathrm{d}\pi(x)\leq d/\alpha$ (c.f. Chewi,, 2026, “basic lemma”). Let $X_{\rho}\sim\rho$, so that $\Psi^{\star}(\|X_{\rho}\|)$ is distributed according to the radial part of $\pi_{\rm rad}^{\star}$. The radial parts of $\rho$ and of $\pi_{\rm rad}^{\star}$ are one-dimensional distributions, with potentials $F_{\rho}$ and $F_{\star}$ respectively; this implies that they are both strongly log-concave, with respective parameters $1$ and $\ell_{V}$. Therefore,
@@ Appendix C Omitted proofs from Section 4
We will closely follow the proof strategy put forth by Jiang et al., (2025) with appropriate modifications throughout.
@@
where $X\sim\rho$. Using standard tools from high-dimensional probability pertaining to Gaussian tail bounds (Van Handel,, 2014), the right-hand side is bounded by $O(\ell_{V}^{-1}\exp(-R^{2}/2))$. Therefore, if $R\gtrsim\sqrt{\log(1/\varepsilon)}$, then $(\mathsf{T1})\leq\varepsilon^{2}/\ell_{V}$.
@@
Next, for $r\leq\sqrt{d}-R$, $\hat{\Psi}(r)=\Psi^{\star}(\sqrt{d}-R)$, and since $\psi^{\star}\leq 1/\sqrt{\ell_{V}}$ everywhere,
$d\int_{0}^{\sqrt{d}-R}(\psi^{\star}-\hat{\psi})^{2}$ $\lesssim d\int_{0}^{\sqrt{d}-R}\bigl(\frac{\Psi^{\star}(\sqrt{d}-R)^{2}}{r^{2}}+\frac{1}{\ell_{V}}\bigr)\,\mathrm{d}\widetilde{\rho}(r)$
$\leq d\Psi^{\star}(\sqrt{d}-R)^{2}\,\mathbb{E}[\left\|X\right\|^{-2}\,\mathbf{1}_{\left\|X\right\|\notin\mathcal{I}}]+\frac{d}{\ell_{V}}\,\mathbb{P}(\left\|X\right\|\notin\mathcal{I})$
$\leq d\Psi^{\star}(\sqrt{d}-R)^{2}\,\mathbb{E}[\left\|X\right\|^{-2.5}]^{4/5}\,\mathbb{P}(\left\|X\right\|\notin\mathcal{I})^{1/5}+\frac{d}{\ell_{V}}\,\mathbb{P}(\left\|X\right\|\notin\mathcal{I})$
$\lesssim d\,\frac{\Psi^{\star}(\sqrt{d}-R)^{2}}{d}\,\mathbb{P}(\left\|X\right\|\notin\mathcal{I})^{1/5}+\frac{d}{\ell_{V}}\,\mathbb{P}(\left\|X\right\|\notin\mathcal{I})$
$\lesssim\frac{d}{\ell_{V}}\exp(-R^{2}/10)\,$,
where we used $d\geq 3$. A similar argument controls the integral over $\sqrt{d}+R,\infty)$. Hence, $R\gtrsim\sqrt{\log(d/\varepsilon)}$ implies $\int_{\|x\|\notin\mathcal{I}}\|D(T^{\star}-\hat{T})(x)\|_{\rm F}^{2}\,\mathrm{d}\rho(x)\lesssim\varepsilon/\ell_{V}$.
@@ C.2 Construction of the Gram matrix
We follow the arguments of Jiang et al., (2025). By Jiang et al., (2025, Propositions 5.10 and 5.11), the objective $\lambda\mapsto\mathcal{F}(T_{\lambda})$ is $\ell_{V}$-strongly convex and $L_{V}\,(1+\Upsilon)$ smooth in the $\|\cdot\|_{Q}$ norm, where $\Upsilon>0$ is the smallest positive constant such that $\|DT_{\lambda}-\alpha I\|_{L^{2}(\rho)}^{2}\leq\Upsilon\,\|T_{\lambda}-\alpha\,{\operatorname{id}}\|_{L^{2}(\rho)}^{2}$ for all $\lambda\in\mathbb{R}^{J+1}$. Writing this out, we need
@@
Let us start with the case of $k\geq 1$. In this case, $r\gtrsim\sqrt{d}$, so the term $\mathsf{T}_{2}$ is bounded, up to an absolute constant, by the integrand on the right-hand side. Thus, for term $\mathsf{T}_{1}$, it suffices to prove $\int_{a_{k}}^{a_{k}+\delta}\,\mathrm{d}\widetilde{\rho}(r)\lesssim\Upsilon\inf_{\bar{r}\in\mathbb{R}}\int_{a_{k}}^{a_{k}+\delta}(r-\bar{r})^{2}\,\mathrm{d}\widetilde{\rho}(r)$, i.e., we need a lower bound on the variance of the distribution $\widetilde{\rho}$ restricted to the interval $a_{k},a_{k}+\delta]$. Note that $\widetilde{\rho}\propto\exp(-F_{\rho})$ where $F_{\rho}$ is defined in the proof of Theorem 3.5. Using the estimates from that proof, we know that $F_{\rho}$ is $O(R)$-Lipschitz on the interval $a_{k},a_{k}+\delta]$, hence $\log\widetilde{\rho}$ only varies by $O(1)$ on this interval provided $\delta\lesssim 1/R$. The argument of Jiang et al., (2025, Lemma 5.13) now shows that the variance is lower bounded by $\Omega(\delta^{2})$, and hence our desired estimate holds for $\Upsilon\asymp\delta^{-2}$.
@@
Recall that for $\lambda\mapsto\mathcal{F}(T_{\lambda})$, the smoothness constant is $\widetilde{O}(J^{2}L_{V})$. Choosing $h=1/\widetilde{\Theta}(L_{V}J^{2})$, we have by standard arguments for projected gradient descent for smooth and strongly convex functions (Beck,, 2017, Theorem 10.29)
@@ C.5 Gradient of the objective
$DT_{\lambda}(x)=\frac{\alpha r+\sum_{j=0}^{J}\lambda_{j}\Psi_{j}(r)}{r}\,I+\Bigl(\sum_{j=0}^{J}\lambda_{j}\Psi_{j}^{\prime}(r)-\frac{1}{r}\sum_{j=0}^{J}\lambda_{j}\Psi_{j}(r)\Bigr)\,\frac{xx^{\top}}{r^{2}}\,$,
@@
$DT_{\lambda}(x)=\Bigl(\alpha+\sum_{j=0}^{J}\lambda_{j}\Psi^{\prime}_{j}(r)\Bigr)\,xx^{\top}/r^{2}+\Bigl(\alpha+\sum_{j=0}^{J}{\lambda_{j}}\Psi_{j}(r)/r\Bigr)\,(I-xx^{\top}/r^{2})\,.$
@@
$\log\det DT_{\lambda}(x)=(d-1)\log\bigl(\alpha+\langle\lambda,\vec{\Psi}(r)/r\rangle\bigr)+\log\bigl(\alpha+\langle\lambda,\vec{\Psi}^{\prime}(r)\rangle\bigr)\,.$
@@
$\nabla_{\lambda_{j}}\,\mathbb{E}_{X\sim\rho}[\log\det DT_{\lambda}(X)]=\int_{0}^{\infty}\Bigl[\frac{(d-1)\,\Psi_{j}(r)}{g_{\lambda}(r)}+\frac{\Psi^{\prime}_{j}(r)}{g^{\prime}_{\lambda}(r)}\Bigr]\,\mathrm{d}\widetilde{\rho}(r)$ (25)
@@
To prove the main result, we follow the strategy of Jiang et al., (2025, Lemma 5.16). Setting up notation, recall that
@@
where we bound $\langle Q^{-1},\vec{\Psi}(\|x\|)\vec{\Psi}(\|x\|)^{\top}\rangle\lesssim J^{3}$. This follows by mimicking Lemma 5.15 by Jiang et al., (2025) but using the computations from the proof of Proposition 4.4; we omit this computation. We bound this last term as follows
@@
The statement follows by employing Theorem 4.3 by Jiang et al., (2025).
@@
We conclude by invoking a standard fact about log-smooth measures (Chewi,, 2026, “basic lemma”). ∎
@@
Our goal is to show that Q{Q} is lower bounded by some absolute constant $c_{Q}>0$. This is sufficient, as
$\langle Q^{-1},\vec{\Psi}(x)\vec{\Psi}(x)^{\top}\rangle{\leq}\frac{1}{c_{Q}}\operatorname{tr}(\vec{\Psi}(x)\vec{\Psi}(x)^{\top})=\frac{1}{c_{Q}}\sum_{j=0}^{J}\Psi_{j}^{2}(\|x\|)\leq\frac{M}{c_{Q}}\,$,
@@ D.1 Hyperparameters
To obtain a Laplace approximation of a posterior, we use the “minimize” function in Scipy (Virtanen et al.,, 2020) with the BFGS optimizer. Our implementation of Forward-Backward VI is exactly as in the publicly available repository by Diao et al., (2023), and we use the same learning rate and number of iterations as our approach.
@@ D.2.1 Student-$t$ distribution
Let $\pi_{\rm Stu}$ be the Student-$t$ distribution with $\nu>0$ degrees of freedom in $\mathbb{R}^{d}$, mean $\mu\in\mathbb{R}^{d}$, and scale matrix $\Sigma\in\mathbb{R}^{d\times d}$. We say $X\sim\pi_{\rm Stu}$ if $X=\mu+Z/\sqrt{W/\nu}$ where $Z\sim\mathcal{N}(0,\Sigma)$ and $W\sim\chi^{2}_{\nu}$ are independent. The density is given by
$\pi_{\rm Stu}(x)=\frac{\Gamma(\frac{\nu+d}{2})}{\Gamma(\frac{\nu}{2})\,(\nu\pi)^{d/2}\,|\Sigma|^{1/2}}\,\Bigl(1+\frac{r(x)^{2}}{\nu}\Bigr)^{-\frac{\nu+d}{2}}\,.$
@@
In the radial setting, it suffices to enforce this on balls of radius $r$ i.e., on sets of the form $\{x\in\mathbb{R}^{d}:\|x\|\leq r\}$, yielding, for a random variable $X\sim\rho$ and $Y\sim\pi,$
@@
and $\Psi^{\star}(r)$} is determined implicitly by solving
@@
If $Y\sim\pi_{\rm Stu}$, then $\|Y\|^{2}/d\sim F_{d,\nu}$, where $F_{d,\nu}$ is the CDF of the $F$-distribution with $(d,\nu)$ degrees of freedom. Hence,
$\mathbb{P}(\|Y\|\leq s)=\mathbb{P}\Bigl(\frac{\|Y\|^{2}}{d}\leq\frac{s^{2}}{d}\Bigr)=F_{d,\nu}\Bigl(\frac{s^{2}}{d}\Bigr)\,.$
@@
$F_{\chi^{2}_{d}}(r^{2})=F_{d,\nu}\Bigl(\frac{\Psi^{\star}(r)^{2}}{d}\Bigr)\,$,
@@
$\Psi^{\star}(r)=\sqrt{d\,F^{-1}_{d,\nu}\bigl(F_{\chi^{2}_{d}}(r^{2})\bigr)}\,.$
@@ D.2.2 Laplace distribution
Solving the implicit equation for the mapped radius $\Psi^{\star}(r)$ point-wise is computationally prohibitive. To circumvent this, we use an interpolation-based approach: the radial CDF of the target is pre-computed on a dense grid, and a linear interpolator of its inverse is constructed. The map then reduces to a direct, vectorized evaluation of this interpolator. Alternatively, the equation can be solved for each point using a root-finding algorithm, such as Brent’s method (Brent,, 1973), although this approach is less efficient for large datasets.